OpenStudy (anonymous):

please help me to integrate this one :

OpenStudy (anonymous):

\[\int\limits \frac{x^4+3}{x^2+1} dx\]

OpenStudy (anonymous):

hint x^4 +3 = x^4 -1 +4

OpenStudy (anonymous):

(x^4 +3 )/( x^2+1) = (x^4 -1 +4 )/( x^2+1) = (( x^2+1)*( x^2-1)) + 4 )/( x^2+1) = ( x^2-1) + 4/( x^2+1)

OpenStudy (anonymous):

i hope u can do further....

OpenStudy (anonymous):

yes i think so ... thanks for your help. @matricked

OpenStudy (anonymous):

yw